# Methodology, Software, Validation, Writing-Original Draft, Writing-Review and Editing, Visualization

### v1

@@ 2.3.3 The whole-communities scale and classification
For the final classification of test subjects, we calculated the Euclidean distance between the feature vectors of the test samples and the two prototypes. Each test sample was then assigned to the category of the prototype with minimal distance, as shown in Figure 3.
@@ 2.4.7 Hi-GCN
Hi-GCN takes the features of each subject extracted by f-GCN as nodes, and gets the edges between subjects through the calculation of correlation coefficients. After that, the population graph networks is bulit. Then, four-layer residual graph convolution is used to achieve feature transfer between subjects, and the learned feature representation of each subject is obtained to realize the final classification task (Jiang et al., 2020).
@@ 6 Loss CrossEntropy
For the hyperparameters $N_{gra}$ and $k$, we set them to be 2 and 7 during the experiments, respectively, because Gong et al. found the first two gradients account for most of the variance of the functional connectivity network (Gong and Zuo, 2023), and Yeo et al. found network template communities (Yeo et al., 2011). In subsection 3.3, we will disscuss the influence of different settings on the classification performance.
@@ 3.1 Performance comparison among different methods
The classification results of the ASD-HNet and compared methods are shown in Table 4, with the same experimental settings. All the results of the compared methods were obtained by reproducing the models according to the public codes provided in the references. The hyper-parameters were set according to the original studies, except for f-GCN and Hi-GCN. During training the f-GCN and Hi-GCN, we set the corresponding optimal threshold to be 0.1 and 0.2 respectively. It can be found that the proposed ASD-HNet outperforms other methods on all four indicators, with accuracy improved by 3.09%, sensitivity improved by 2.94%, specificity improved by 1.38% and F1-score improved by 2.45%, respectively.
@@
Method Accuracy(%) Sensitivity(%) Specificity(%) F1-score(%)
f-GCN* 61.32 61.28 61.31 59.13
BrainnetCNN* 64.52 59.84 68.46 61.03
CNN-model* 64.88 58.55 69.70 60.36
FBNETGEN* 67.14 69.69 64.65 68.59
ASD-DiagNet* 66.70 60.64 71.91 62.56
BrainTransformerNet* 67.00 64.23 70.23 67.14
Hi-GCN* 67.34 68.78 65.70 65.17
MCG-Net* 70.48 67.08 73.52 68.32
ASD-HNet 73.57 72.63 74.90 71.04
@@
For a intuitive comparison between these method and our ASD-HNet, we directly quoted the reported results from origninal references as some previous studies (Wang et al., 2022). In these studies, the authors did not release the implementation codes, or did not describe how to generate the used data from ABIDE dataset. The comparison results are summarized in Table 5. By this rough comparison, we can also find that although the F1-score of our method is little lower than MBT and DGAIB methods, our method achieved the best accuracy (73.57%), sensitivity (72.63%) and specificity (74.90%).
@@ 3.2 Ablation experiment
To analyze the rationality and effectiveness of the architecture of the proposed model, we conducted several ablation experiments to check whether removing each key module will degrade the overall performance. Accordingly, in order to verify that the ROI-scale graph convolution has effective information transfer and whether the ROI-attention module is effective in improving the model performance, we conducted a comparative test with/without GCN and ROI-attention. Similarly, we only used the orignal linear layer during training the feature extractor to achieve the classification to test the validity of prototype learning. The results of the ablation experiment were shown in Table 6.
@@ 3.3 The influence of the hyperparameters settings on the model performance
In ASD-HNet, two hyperparameters $N_{gra}$ and $k$ need to be configured, which represent the number of functional gradients components and the number of categories for clustering, respectively. The results reported above were obtained when $N_{gra}$ and $k$ were set to be 2 and 7. Different combination of these two hyperparameters may yield different classification performance. Therefore, we carried out experiments to check the influence of the hyperparameters settings on the model performance, with $k$ ranged from 5 to 10 and $N_{gra}$ ranged from 2 to 6. The experimental results are shown in Figure 4. The highest score for each indicator is marked with a red circle.
From the results in the Figure 4, we can find that although the best combination of hyperparameter values in the sensitivity index is different from the other three indicators, the three indicators of accuracy, specificity and F1-score all achieve the best scores when $k$=7 and $N_{gra}$=2. Moreover, sensitivity has the third highest score under this combination of hyperparameter values. Based on comprehensive evaluation, setting the $k$=7 and $N_{gra}$ as 7 and 2 was a viable choice. Besides, these settings matched the findings in previous studies, such as the first two gradients can account for most of the variance of the functional connectivity network (Gong and Zuo, 2023), and there exists Yeo 7 network template communities (Yeo et al., 2011).
Figure 4: The averaged four metrics obtained by different combination of $k$ and $N_{gra}$.
@@ 3.4 The interpretability and visualization analysis
In this section, we investigate whether the proposed model can find the discriminative features that can be explained with previous neural mechanisms.
First, we want to check whether the model can find the difference patterns of FCN between ASD subjects and HCs subjects. F-score is a technique for extracting key features based on a data-driven approach. For the analysis of key edges of brain regions, we retained the $50\%$ edges with the highest F-score for retention and visualization, as shown in Figure 5. All the visualization are achieved with BrainNet Viewer (Xia et al., 2013) tool. The most differentiated edges observed in Figure 5 are mostly the connections between the temporal lobe, the frontal lobe and the sensorimotor area, which is consistent with previous findings. For instance, Abrams et al. has reported an atypical decline in functional connectivity between the temporal area (auditory network) and the medial prefrontal area (DMN) in people with autism (Abrams et al., 2013).
@@
Furthermore, we want to show that the functional gradients in the model can find the difference patterns on the community networks. We conducted the analysis of the validity and interpretability of utilizing the first two components of functional gradients for constructing community networks. Since the gradients calculated separately for the two groups of levels may not be directly comparable (Vos de Wael et al., 2020), we chose the joint embedding (Xu et al., 2020) to realize the alignment of the two groups of functional gradients. Then we conducted individual analyses of both the principal and secondary gradients, as well as a joint analysis of the principal and secondary gradients.
The aligned functional gradients of the ASD and HC groups are used to find the ten ROIs with the largest difference in absolute values of the principal gradients of the two groups. The same method is used to find the ten ROIs with the biggest difference in the secondary gradient between the two groups. The brain regions with the greatest differences in the principal gradients are found to be right parahippocampal gyrus (PHG.R), left superior frontal gyrus of orbital part (ORBsup.L), right superior parietal gyrus (SPG.R), right cerebellum inferior (CRBLCrus2.R), left middle occipital gyrus (MOG.L), left middle frontal gyrus (MFG.L), vermis (Vermis45), right precuneus (PCUN.R), right paracentral lobule (PCL.R), right hippocampus (HIP.R). The result is shown in the Figure 6.
The ten ROIs with the largest difference in the secondary gradient are right Supplementary motor area (SMA.R), left Precental gyrus (PreCG.L), left Posterior cingulate gyrus (PCG.L), right Inferior parietal, but supramarginal and angular gyri (IPL.R), left Middle occipital gyrus (MOG.L), Vermis (Vermis45), right Inferior frontal gyrus of orbital part (ORBinf.R), left Middle temporal gyrus (MTG.L), right Precuneus (PCUN.R), right Middle temporal gyrus (MTG.R). The result is shown in the Figure 7.
Figure 6: The ten ROIs with the largest difference in group principal gradients.
Figure 7: The ten ROIs with the largest difference in the secondary gradients.
@@
Right after that, we merged the first two gradients (principal and secondary gradients) and employed Euclidean distance (consistent with K-Means algorithm) to measure the distinction in corresponding ROIs between the ASD and HC groups. Subsequently, we visualized the ten brain regions with the greatest Euclidean distance (indicating the most significant difference). The visualized result is shown in Figure 8.
@@
Figure 8: The ten ROIs with the greatest differences in the first two gradients.
Based on the aforementioned findings, disparities in the attention towards specific brain regions are observed between the principal and secondary gradients. For instance, the principal gradient demonstrated a higher focus on the left frontal lobe, whereas the secondary gradient prioritizes the left temporal lobe. By concurrently considering the first two gradients, a more comprehensive evaluation of functional differences and similarities among brain regions can be attained, consequently leading to an improved community clustering effect. Our conclusion suggested that the variations in brain regions among ASD patients primarily manifest in the superior frontal gyrus of the orbital part, precuneus, and several regions within the temporal lobe. This is similar to (Hong et al., 2019) in discovery roughly.
@@ 3.5 Experiments on model generalization
In order to investigate the generalization of our proposed method, as some previous studies (Zhang et al., 2023; Liu et al., 2023b), we used the proposed model to conduct classification on a ADHD-200 Consortium (ADHD-200) dataset  [^1] http://fcon_1000.projects.nitrc.org/indi/adhd200/. We choose the dataset preprocessed by Athena pipeline (Bellec et al., 2017). A total of 759 subjects from 7 sites were contained for fMRI data, including 303 ADHD patients and 456 HCs. Similar to the above experiments on ASD, we used 10-fold cross-validation to conduct corresponding experiments. Initially, for ASD-HNet, we used the same hyperparameters as we did for classifying ASD. We then tuned the model based on the specifics of the ADHD data. The adjustments were as follows: (1) Due to the imbalance in the ADHD data, we used a larger batchsize (64) to ensure that each batch included a more balanced representation of both classes. (2) Given the presence of multiple subtypes of ADHD (Molavi et al., 2020), we incorporated more global brain features from different perspectives to enhance the model’s discriminative ability. Specifically, we increased the out__\__channels of conv2d in the global brain scale to 32. These adjustments can enhance the model’s adaptation and optimization for classifying ADHD.
In addition to the comparative methods mentioned above for ASD, we also included two state-of-the-art methods in our analysis: (1) MDCN (Yang et al., 2023), which utilize fMRI data from 572 subjects, including 261 ADHD patients and 311 healthy controls (HCs); (2) A-GCL (Zhang et al., 2023), which uses data from 947 subjects, comprising 362 ADHD patients and 585 HCs. The experimental results are shown in Table 7. The results indicate that our method also achieves better performance in classifying ADHD. Due to the imbalance in the data we used, the comparison methods we reproduced show a significant imbalance in specificity and sensitivity, resulting in a lower F1-score. However, our method achieves the best performance in terms of accuracy (71.33%) and specificity (80.30%). This demonstrated that ASD-HNet not only performs well in classifying ASD but also achieved good classification performance for ADHD, which validates its robustness and generalization.
@@ 3.6 Limitation
Even though ASD-HNet achieved better performance than other methods, some limitations still exist in this study. First of all, the classification performance is still unsatisfactory and can not be applied to clinical scenes in current stage. External cohort dataset should be collected for further validation (Han et al., 2019; Wang et al., 2021). The proposed method relies on a large amount of data for corresponding training. When the amount of data is insufficient or unbalanced, how to maintain the performance of the model is an important issue worthy of further research. Besides, we did not consider the large variability among the dataset from different sites into the model construction. It is beneficial to explicitly eliminate the differences on data distribution with some technologies, such as domain adaption (Wang et al., 2019; Wen et al., 2024; Liu et al., 2023c). In future, we need address these issues during designing models to conduct the multi-site fMRI data analysis.

### v2

@@ 2.3.3 The global representation scale and classification
For the final classification of test subjects, we calculated the Euclidean distance between the feature vectors of the test samples and the two prototypes. Each test sample was then assigned to the class of the prototype with minimal distance, as shown in Figure 3.
@@ 2.4.7 Hi-GCN
Hi-GCN takes the features of each subject extracted by f-GCN as nodes, and gets the edges between subjects through the calculation of correlation coefficients. After that, the population graph networks is built. Then, four-layer residual graph convolution is used to achieve feature transfer between subjects, and the learned feature representation of each subject is obtained to realize the final classification task (Jiang et al., 2020).
@@ 6 Loss CrossEntropy
For the hyperparameters $N_{gra}$ and $k$, we set them to be 2 and 7 during the experiments, respectively. In subsection 3.3.1, we will discuss the influence of different settings on the classification performance.
@@ 3.1 Performance comparison among different methods
The classification results of the ASD-HNet and compared methods are shown in Table 4, with the same experimental settings. All the results of the compared methods were obtained by reproducing the models according to the public codes provided in the references. The hyper-parameters were set according to the original studies, except for f-GCN and Hi-GCN. During training the f-GCN and Hi-GCN, we set the corresponding optimal threshold to be 0.1 and 0.2 respectively. It can be found that the proposed ASD-HNet outperforms other methods on all four indicators, with accuracy improved by 3.09%, sensitivity improved by 2.94%, specificity improved by 1.38% and F1-score improved by 2.45%, respectively. To further validate the effectiveness of the proposed method, we performed a paired t-test on the ten-fold cross-validation accuracy of the models. The results reveal that the p-values for comparisons between ASD-HNet and all other methods are less than 0.05, demonstrating the statistical significance of the performance improvements achieved by the proposed method.
@@
Method Accuracy(%) Sensitivity(%) Specificity(%) F1-score(%) p-value
f-GCN* 61.32 61.28 61.31 59.13 0.0002
BrainnetCNN* 64.52 59.84 68.46 61.03 0.0006
CNN-model* 64.88 58.55 69.70 60.36 0.0053
FBNETGEN* 67.14 69.69 64.65 68.59 0.0029
ASD-DiagNet* 66.70 60.64 71.91 62.56 0.0023
BrainTransformerNet* 67.00 64.23 70.23 67.14 0.0151
Hi-GCN* 67.34 68.78 65.70 65.17 0.0052
MCG-Net* 70.48 67.08 73.52 68.32 0.0375
ASD-HNet 73.57 72.63 74.90 71.04 -
@@
For a intuitive comparison between these method and our ASD-HNet, we directly quoted the reported results from original references as previous studies (Wang et al., 2022). In these studies, the authors did not release the implementation codes, or did not describe how to generate the used data from ABIDE dataset. The comparison results are summarized in Table 5. By this rough comparison, we can also find that although the F1-score of our method is little lower than MBT and DGAIB methods, our method achieved the best accuracy (73.57%), sensitivity (72.63%) and specificity (74.90%).
@@ 3.2 Ablation experiment
To evaluate the rationality and effectiveness of the proposed model’s architecture, we conducted a series of ablation experiments to assess the impact of removing each key module on overall performance. Specifically, to validate the effectiveness of information transfer in the ROI-scale graph convolution and to determine whether the ROIs attention module contributes to improved model performance, we performed comparative experiments with and without the graph convolutional network and the ROIs attention mechanism. Additionally, to test the validity of prototype learning, we trained the feature extractor using only the original linear layer for classification, thereby isolating the contribution of prototype learning to the model’s performance. The results of the ablation experiment were shown in Table 6.
@@ 3.3 The influence of the hyperparameters settings on the model performance
3.3.1 Influence of the $N_{gra}$ and $k$
In ASD-HNet, two hyperparameters $N_{gra}$ and $k$ need to be configured, which represent the number of functional gradients components and the number of functional communities, respectively. The results reported above were obtained when $N_{gra}$ and $k$ were set to be 2 and 7. Different combinations of these two hyperparameters may lead to varying classification performance. To investigate the impact of hyperparameter settings on model performance, we conducted the systematic hyperparameter search experiments as previous studies (Jiang et al., 2020; Wen et al., 2022; Wang et al., 2024). We conducted experiments with $k$ ranging from 5 to 10 and $N_{gra}$ ranging from 2 to 6. The experimental results are illustrated in Figure 4. The highest score for each metric is marked with a red circle.
From the results in the Figure 4, we can find that although the best combination of hyperparameter values in the sensitivity index is different from the other three indicators, the three indicators of accuracy, specificity and F1-score all achieve the best performance when $k$=7 and $N_{gra}$=2. Moreover, sensitivity has the third highest score under this combination of hyperparameter values. In previous study, Gong et al. found that the first two functional gradients can explain the majority of variance in functional networks (Gong and Zuo, 2023). Similarly, Samara et al. also found similar results in their research (Samara et al., 2023). Our experimental results are consistent with previous findings. Based on comprehensive evaluation, we set the $k$ and $N_{gra}$ as 7 and 2 in all our experiments.
Figure 4: The averaged four metrics obtained by different combination of the parameters $k$ and $N_{gra}$.
3.3.2 Influence of the Node Feature Retention Ratio
In the ROIs scale, there exists a hyperparameter ${\lambda_{R}}$ (node feature retention ratio). To investigate the impact of this parameter on the overall structural performance, we conducted a hyperparameter search experiment. The ${\lambda_{R}}$ ranges from 30% to 60% with a step size of 5%. The experimental results are presented in Figure 5. It can be found that when the retention ratio is set at 50%, the model achieves optimal performance in terms of accuracy, sensitivity, and F1-score, while the specificity metric is only slightly lower than that at the 60% ratio. An appropriate retention ratio enables the model to better learn useful features. When the retention ratio is too low, some features beneficial for classification may be inadvertently removed. Conversely, when the retention ratio is too high, an excessive number of noisy features may be retained. Both scenarios can lead to suboptimal model performance. Therefore, we set ${\lambda_{R}}$ to be 50% in current study to report the experimental results.
Figure 5: Impact of node feature retention ratio (${\lambda_{R}}$) on model performance.
@@ 3.4 The interpretability and visualization analysis
In this section, we explore the interpretability of the proposed method and the learned ROIs attention, investigating whether the discriminative features identified are interpretable and consistent with previous findings.
First, we want to check whether the model can find the difference patterns of FCN between ASD subjects and HCs subjects. F-score is a technique for extracting key features based on a data-driven approach. For the analysis of key edges of brain regions, we retained the $10\%$ edges with the highest F-score for retention and visualization, as shown in Figure 6. All the visualization are achieved with BrainNet Viewer (Xia et al., 2013) tool. The most differentiated edges observed in Figure 6 are mostly the connections between the temporal lobe, the frontal lobe and the sensorimotor area, which is consistent with previous findings. For instance, Abrams et al. has reported an atypical decline in functional connectivity between the temporal area (auditory network) and the medial prefrontal area (DMN) in people with autism (Abrams et al., 2013).
@@
Furthermore, we want to show that the functional gradients in the model can find the difference patterns on the community networks. We conducted the analysis of the validity and interpretability of utilizing the first two components of functional gradients for constructing community networks. Since the gradients calculated separately for the two groups of levels may not be directly comparable (Vos de Wael et al., 2020), we chose the joint embedding (Xu et al., 2020b) to realize the alignment of the two groups of functional gradients. Then we conducted a joint analysis of the principal and secondary gradients.
@@
We merged the first two gradients (principal and secondary gradients) and employed Euclidean distance (consistent with K-Means algorithm) to measure the distinction in corresponding ROIs between the ASD and HC groups. Subsequently, we visualized the ten brain regions with the greatest Euclidean distance (indicating the most significant difference). The visualized result is shown in Figure 7(a).
(a) The ten ROIs with the greatest differences in the first two gradients.
(b) The ten ROIs with the highest attention scores obtained from ROIs attention scores.
@@
Figure 7: The ten ROIs with the greatest differences in the first two functional gradients and the ten ROIs with the most significant differences learned by ROIs attention weights.
Based on the aforementioned findings, we can observe that the group-level functional gradient differences are primarily reflected in the following brain regions: the inferior orbital gyrus (ORBinf.R), right cerebellar Crus II (CRBLCrus2.R), left middle temporal gyrus (MTG.L), left middle frontal gyrus (MFG.L), right hippocampus (HIP.R), right precuneus (PCL.R), right middle temporal gyrus (MTG.R), left middle occipital gyrus (MOG.L), vermis (Vermis45), and right precuneus (PCUN.R). The MTG has been widely confirmed to be associated with ASD (Xu et al., 2020a). And the cerebellum, due to its crucial role in regulating brain function, has significant associations with ASD (Mosconi et al., 2015; Hampson and Blatt, 2015). Xiao et al. also confirmed the association between the precuneus and ASD (Xiao et al., 2023). These key ROIs and their associations are consistent with the findings of (Hong et al., 2019).
Next, we visualize the ROIs attention weights learned by the model. The ten ROIs with the highest attention scores are visualized in Figure 7(b). These ROIs include: left cerebellum (Cere3.L, Cere10.L), left dorsal cingulate gyrus (DCG.L), right olfactory cortex (OLF.R), right triangular part of the inferior frontal gyrus (IFGtriang.R), right cerebellum (Cere8.R), right middle orbital gyrus (ORBmid.R), vermis6, right middle temporal gyrus (MTG.R), and left middle frontal gyrus (MFG.L). We observe that several regions overlap between the group-level differences identified by the functional gradient and those highlighted by the ROIs attention scores, particularly in the temporal lobe (Xu et al., 2020a), left prefrontal cortex (Lee et al., 2009), and certain cerebellar areas (Mosconi et al., 2015; Hampson and Blatt, 2015).
@@ 3.5 Experiments on model generalization
In order to investigate the generalization performance of the proposed method, we conducted experiments on both the ADHD-200 Consortium (ADHD-200) dataset  [^1] http://fcon_1000.projects.nitrc.org/indi/adhd200/ and the ABIDE-II dataset (Di Martino et al., 2017).
3.5.1 ADHD-200 Dataset
ADHD-200 dataset has 973 resting state fMRI acquisitions labeled with ADHD or healthy controls from subjects aging from 7 to 21. The dataset is collected from eight different international sites, which are NeuroImage group (NeuroImage), New York University Child Study Center (NYU), Peking University (Peking), Brown University (BU), Kennedy Krieger Institute (KKI), University of Pittsburgh (Pitts), Oregon Health and Science University (OHSU) and Washington University in St. Louis (WU).
We choose the dataset preprocessed by Athena pipeline (Bellec et al., 2017). A total of 759 subjects from 7 sites were contained for fMRI data, including 303 ADHD patients and 456 HCs. Similar to the above experiments on ASD, we used 10-fold cross-validation to conduct corresponding experiments. Initially, for ASD-HNet, we used the same hyperparameters as we did for classifying ASD. We then tuned the model based on the specifics of the ADHD data. The adjustments were as follows: (1) Due to the imbalance in the ADHD data, we used a larger batchsize (64) to ensure that each batch included a more balanced representation of both classes. (2) Given the presence of multiple subtypes of ADHD (Molavi et al., 2020), we incorporated more global brain features from different perspectives to enhance the model’s discriminative ability. Specifically, we increased the out__\__channels of Conv2d in the global brain scale to 32. These adjustments can enhance the model’s adaptation and optimization for classifying ADHD.
In addition to the comparative methods mentioned above for ASD, we also included two state-of-the-art methods in our analysis: (1) MDCN (Yang et al., 2023), which utilize fMRI data from 572 subjects, including 261 ADHD patients and 311 healthy controls (HCs); (2) A-GCL (Zhang et al., 2023b), which uses data from 947 subjects, comprising 362 ADHD patients and 585 HCs. The experimental results are shown in Table 7. The results indicate that our method also achieves better performance in classifying ADHD. Due to the imbalance in the data we used, the comparison methods we reproduced show a significant imbalance in specificity and sensitivity, resulting in a lower F1-score. However, our method achieves the best performance in terms of accuracy (71.33%) and specificity (80.30%). This demonstrated that ASD-HNet not only performs well in classifying ASD but also achieved good classification performance for ADHD, which validates its robustness and generalization.
@@ 3.5.1 ADHD-200 Dataset
3.5.2 ABIDE-II Dataset
The ABIDE-II initiative was established as an extension of ABIDE-I, aiming to further advance scientific research on the brain connectome in Autism Spectrum Disorder. ABIDE-II has aggregated over 1000 additional datasets, offering more comprehensive phenotypic characterizations, particularly in the quantitative assessment of core ASD symptoms and associated features. The project now encompasses 19 research sites, including 10 founding institutions and 7 newly participating members, collectively contributing 1114 datasets from 521 individuals with ASD and 593 healthy controls (age range: 5–64 years). These data were openly released to the global scientific community in June 2016, providing a critical resource for ASD research.
In addition to the ADHD-200 dataset, we also evaluated the generalization performance of ASD-HNet on the ABIDE-II dataset, which consists of fMRI data from individuals with autism spectrum disorder and matched healthy controls. For this dataset, we used data from five sites, i.e., BN, EMC, GU, NYU, and OHSU, including a total of 382 subjects, comprising 185 ASD patients and 197 healthy controls. The preprocessing procedure for this dataset followed the same pipeline as ABIDE-I. Since the sample size of this dataset is smaller than that of the ABIDE-I and ADHD-200 datasets, we employed 5-fold cross-validation to balance training efficiency and validation accuracy. The experimental results are presented in Table 8, five compared methods were selected as the baselines.
The experimental results demonstrate that the ASD-HNet method achieves the best performance, which achieves performance improvements of 4.43% in accuracy, 7.47% in specificity, and 2.32% in F1-score compared to the the second-best method. These additional results further validating the generalization ability of the proposed method across different datasets for the same disease.
Table 8: Experimental results of ASD-HNet and compared methods on ABIDE-II data, with the best results are highlighted in bold.
ABIDE-II Method ACC(%) SEN(%) SPE(%) F1-score(%)
f-GCN 62.40 56.20 60.30 52.33
BrainnetCNN 68.12 69.72 62.12 66.43
CNN-model 66.85 59.21 70.57 62.49
Hi-GCN 64.17 64.17 56.58 56.63
MCG-Net 65.51 61.55 64.06 62.00
ASD-HNet(ours) 72.55 65.23 78.04 68.75
@@ 3.6 Limitation
Even though ASD-HNet achieved better performance than other methods, some limitations still exist in this study. First of all, the classification performance is still unsatisfactory and can not be applied to clinical scenes in current stage. External cohort datasets should be collected for further validation (Han et al., 2019; Wang et al., 2021). The proposed method relies on a large amount of data for corresponding training. When the amount of data is insufficient or unbalanced, how to maintain the performance of the model is an important issue worthy of further research. Besides, we did not consider the large variability among the dataset from different sites into the model construction in current study. It is beneficial to explicitly eliminate the differences on data distribution with some technologies, such as domain adaption (Wang et al., 2019; Wen et al., 2024; Liu et al., 2023b). We have also recognized the issue of site heterogeneity and have conducted preliminary experiments to mitigate its impact, as discussed in (Luo et al., 2025). In the future, we need address these issues during designing models to conduct the multi-site fMRI data analysis.